OpenStudy (anonymous):

Solve the equation 4n - 3 = 2n + 5 for n.

OpenStudy (anonymous):

4n-3=2n+5 +3 +3 4n=2n+8 -2n -2n 2n=8 divide each side by 2 n=4

OpenStudy (anonymous):

4n-3=2n+5, 4n-2n-3=5, 2n-3=5, 2n=5+3, 2n=8, n=8/2, n=4.